OpenStudy (anonymous):

Confused m+3/4=3/4m+3/2 can someone help?i know I have to find the lcm but idk how?

OpenStudy (anonymous):

do you mean solving the equation? what is lcm? lcm for 2 numbers? m+3/4 = 3/4m + 3/2 m - 3/4m = 3/2 - 3/4 1/4m = 3/4 m = 3

OpenStudy (anonymous):

yeah I don't know you solve it but thx